A jet lands on a aircraft carrier at 140 mile (63 m/s)
(a)what is its acceleration if it stops in 2 second?
(b)what is the displacement of the plane.which it is stopping?
1
Expert's answer
2020-01-31T10:44:16-0500

a)


"a=\frac{v}{t}"

"a=\frac{63}{2}=31.5\frac{m}{s^2}"

b)


"s=0.5at^2"

"s=0.5(31.5)(2)^2=63\ m"